Support Department Cost Allocation—Direct Method

Yo-Down Inc. produces yogurt. Information related to the company’s yogurt production follows:

Production
Department 1
Production
Department 2
Production
Department 3
Support Department 1 cost driver 1,400       100       500      

Support Department 1’s costs total $142,000. Using the direct method of support department cost allocation, determine the costs from Support Department 1 that should be allocated to each production department.

Production
Department 1
Production
Department 2
Production
Department 3
Support Department 1 cost allocation
Answer #1

Support department total cost drivers = Production Department 1 + Production Department 2 + Production Department 3

= 1,400+100+500

= 2,000

Production
Department 1
Production
Department 2
Production
Department 3
Support Department 1 allocation 142,000 x 1,400/2,000 = $99,400 142,000 x 100/2,000 = $7,100 142,000 x 500/2,000 = $35,500

Kindly comment if you need further assistance. Thanks‼!